OpenStudy (anonymous):

x + 5x + 3x + 30 = 180 x=? A)x=16.67 B)x=18 C)x=24

OpenStudy (anonymous):

JUST SOLVE

OpenStudy (unklerhaukus):

\[x + 5x + 3x + 30 = 180\] \[(1+5+3)x+ 30 = 180\]\[(1+5+3)x = 180-30\]\[x = \frac{180-30}{(1+5+3)}\]

OpenStudy (jiteshmeghwal9):

combine like terms first x+5x+3x+30=180 9x+30=180 9x=180-30 9x=150 x=150/9=?

OpenStudy (anonymous):